# Unconditional stability of a recurrent neural circuit implementing divisive normalization

### v2

@@ 3 Model description
$\begin{split}\boldsymbol{\tau}_{y}\odot\dot{\mathbf{y}}&=-\mathbf{y}+\mathbf{b}\odot\mathbf{z}+\left(\mathbf{1}-\mathbf{a}^{+}\right)\odot\mathbf{W}_{r}\left(\sqrt{\mathbf{y}^{+}}-\sqrt{\mathbf{y}^{-}}\right)\\ \boldsymbol{\tau}_{a}\odot\dot{\mathbf{a}}&=-\mathbf{a}+\mathbf{b}_{0}^{2}\odot\boldsymbol{\sigma}^{2}+\mathbf{W}\,\left(\left(\mathbf{y}^{+}+\mathbf{y}^{-}\right)\odot\mathbf{a}^{+2}\right)\end{split}$ (1)
@@ 5 Stability analysis for arbitrary recurrent weights
$\begin{split}\boldsymbol{\tau}_{y}\odot\dot{\mathbf{y}}&=-\mathbf{y}+\mathbf{b}\odot\mathbf{z}+\left(\mathbf{1}-\sqrt{\lfloor\mathbf{a}\rfloor}\right)\odot\mathbf{W}_{r}\mathbf{y}\\ \boldsymbol{\tau}_{a}\odot\dot{\mathbf{a}}&=-\mathbf{a}+\mathbf{b}_{0}^{2}\odot\boldsymbol{\sigma}^{2}+\mathbf{W}\,\left(\mathbf{y}^{2}\odot\lfloor\mathbf{a}\rfloor\right)\end{split}$ (15)
@@ 7 Discussion
The authors acknowledge valuable discussions with Flaviano Morone, Asit Pal, Mathias Casiulis, and Guanming Zhang. This work was supported by the National Institute of Health under award number R01EY035242. S.M. acknowledges the Simons Center for Computational Physical Chemistry for financial support. This work was supported in part through the NYU IT High Performance Computing resources, services, and staff expertise.
@@ Appendix A Derivation of ORGaNICs
$\begin{split}\boldsymbol{\tau}_{y}\odot\dot{\mathbf{y}}&=-\mathbf{y}+\mathbf{b}\odot\mathbf{z}+\left(\mathbf{1}-\mathbf{a}^{+}\right)\odot\mathbf{W}_{r}\left(\sqrt{\mathbf{y}^{+}}-\sqrt{\mathbf{y}^{-}}\right)\\ \boldsymbol{\tau}_{a}\odot\dot{\mathbf{a}}&=-\mathbf{a}+\mathbf{b}_{0}^{2}\odot\boldsymbol{\sigma}^{2}+\mathbf{W}\,\left(\left(\mathbf{y}^{+}+\mathbf{y}^{-}\right)\odot\mathbf{a}^{+2}\right)\end{split}$ (34)
@@
$\begin{split}\boldsymbol{\tau}_{y}\odot\dot{\mathbf{y}}&=-\mathbf{y}+\mathbf{b}\odot\mathbf{z}+\left(\mathbf{1}-\mathbf{a}^{+}\right)\odot\mathbf{W}_{r}\left(\left(\mathbf{y}^{+}\right)^{1/p}-\left(\mathbf{y}^{-}\right)^{1/p}\right)\\ \boldsymbol{\tau}_{a}\odot\dot{\mathbf{a}}&=-\mathbf{a}+\frac{\mathbf{a}}{\left(\mathbf{a}^{+}\right)^{p}}\odot\left[\mathbf{b}_{0}^{p}\odot\boldsymbol{\sigma}^{p}+\mathbf{W}\left(\left(\mathbf{y}^{+}+\mathbf{y}^{-}\right)\odot\left(\mathbf{a}^{+}\right)^{p}\right)\right]\end{split}$ (35)
@@ Appendix F Analysis of the Rectified model
$\begin{split}\boldsymbol{\tau}_{y}\odot\dot{\mathbf{y}}&=-\mathbf{y}+\mathbf{b}\odot\mathbf{z}+\left(\mathbf{1}-\mathbf{a}^{+}\right)\odot\mathbf{W}_{r}\left(\sqrt{\mathbf{y}^{+}}\right)\\ \boldsymbol{\tau}_{a}\odot\dot{\mathbf{a}}&=-\mathbf{a}+\mathbf{b}_{0}^{2}\odot\boldsymbol{\sigma}^{2}+\mathbf{W}\,\left(\mathbf{y}^{+}\odot\mathbf{a}^{+2}\right)\end{split}$ (87)
@@ F.3 Linear stability analysis of the two-dimensional model
$\begin{split}\tau_{y}\dot{y}&=-y+bz+\left(1-\sqrt{\lfloor a\rfloor}\right)w_{r}\lfloor y\rfloor\\ \tau_{a}\dot{a}&=-a+b_{0}^{2}\sigma^{2}+w\lfloor y\rfloor^{2}\lfloor a\rfloor\end{split}$ (102)
@@ Appendix G Iterative algorithm
$\begin{split}\boldsymbol{\tau}_{y}\odot\dot{\mathbf{y}}&=-\mathbf{y}+\mathbf{b}\odot\mathbf{z}+\left(\mathbf{1}-\sqrt{\lfloor\mathbf{a}\rfloor}\right)\odot\mathbf{W}_{r}\mathbf{y}\\ \boldsymbol{\tau}_{a}\odot\dot{\mathbf{a}}&=-\mathbf{a}+\mathbf{b}_{0}^{2}\odot\boldsymbol{\sigma}^{2}+\mathbf{W}\,\left(\mathbf{y}^{2}\odot\lfloor\mathbf{a}\rfloor\right)\end{split}$ (109)
@@ I.1 Static MNIST input
$\begin{split}\boldsymbol{\tau}_{y}\odot\dot{\mathbf{y}}&=-\mathbf{y}+f(\mathbf{W}_{bx}\mathbf{x})\odot\left(\mathbf{W}_{zx}\mathbf{x}\right)+\left(\mathbf{1}-\sqrt{\lfloor\mathbf{a}\rfloor}\right)\odot\mathbf{W}_{r}\mathbf{y}\\ \boldsymbol{\tau}_{a}\odot\dot{\mathbf{a}}&=-\mathbf{a}+\mathbf{b}_{0}^{2}\odot\boldsymbol{\sigma}^{2}+\mathbf{W}\,\left(\mathbf{y}^{2}\odot\lfloor\mathbf{a}\rfloor\right)\end{split}$ (134)
@@ I.2 Permuted and Unpermuted sequential MNIST
We performed a random split of 57,000 training samples and 3,000 validation samples and picked the model with the largest validation accuracy for testing. The unpermuted sequential MNIST task is defined as follows: for a given $28\times 28$ image, we flatten it to get a one-dimensional, 784 timestep-long input. Then these pixels are presented as an input ($\mathbf{x}_{i}$, one pixel at each time-step $i$) to the discretized ORGaNICs model given by the equation,
$\begin{split}\mathbf{y}_{i+1}&=\mathbf{y}_{i}+\frac{\Delta t}{\boldsymbol{\tau}_{y}}\odot\left(-\mathbf{y}_{i}+\mathbf{b}_{i}\odot\left(\mathbf{W}_{zx}\mathbf{x}_{i}\right)+\left(\mathbf{1}-\mathbf{a}_{i}^{+}\right)\odot\mathbf{W}_{r}\left(\sqrt{\mathbf{y}_{i}^{+}}\right)\right)\\ \mathbf{a}_{i+1}&=\mathbf{a}_{i}+\frac{\Delta t}{\boldsymbol{\tau}_{a}}\odot\left(-\mathbf{a}_{i}+\mathbf{b}_{0,i}^{2}\odot\boldsymbol{\sigma}^{2}+\mathbf{W}\,\left(\mathbf{y}_{i}^{+}\odot\mathbf{a}_{i}^{+2}\right)\right)\\ \mathbf{b}_{i+1}&=\mathbf{b}_{i}+\frac{\Delta t}{\boldsymbol{\tau}_{b}}\odot\left(-\mathbf{b}_{i}+f(\mathbf{W}_{bx}\mathbf{x}_{i}+\mathbf{W}_{by}\mathbf{y}_{i}+\mathbf{W}_{ba}\mathbf{a}_{i})\right)\\ \mathbf{b}_{0,i+1}&=\mathbf{b}_{0,i}+\frac{\Delta t}{\boldsymbol{\tau}_{b_{0}}}\odot\left(-\mathbf{b}_{0,i}+f(\mathbf{W}_{b_{0}x}\mathbf{x}_{i}+\mathbf{W}_{b_{0}y}\mathbf{y}_{i}+\mathbf{W}_{b_{0}a}\mathbf{a}_{i})\right)\end{split}$ (135)
When we are done presenting the pixels we use $\mathbf{y}^{+}_{784}=\lfloor\mathbf{y}_{784}\rfloor^{2}$ to make the predictions. To make this more challenging we also train ORGaNICs on permuted sMNIST where we first permute the pixels of all the images in some random order and the rest of the task is the same. Instead of parametrizing $\boldsymbol{\tau}$, we parametrize ${\Delta t}/{\boldsymbol{\tau}_{y}}=0.05*f(\mathbf{p}_{y})$, ${\Delta t}/{\boldsymbol{\tau}_{a}}=0.01*f(\mathbf{p}_{a})$ and ${\Delta t}/{\boldsymbol{\tau}_{b}}=0.1*f(\mathbf{p}_{b})$ and ${\Delta t}/{\boldsymbol{\tau}_{b_{0}}}=0.1*f(\mathbf{p}_{b_{0}})$, so we can control the dimensionless relative time constants. In practice, we find it is better to make the $\mathbf{a}$ neurons sluggish compared to $\mathbf{y}$. This is based on the intuition given by the two-dimensional phase portrait for different relative time constants Fig. 3. All the parameters (including $\mathbf{W}_{r}$) are unconstrained for this task with initialization specified in Table 6. Since ORGaNICs are stable, we did not need to use gradient clipping for training, which is commonly used for LSTMs. Additionally, we train the model using a StepLR learning rate scheduler with parameters given in Table 7. The model is trained using backpropagation through time (BPTT) and takes approximately 30 hours to fully train.

### v3

@@ 3 Model description
$\begin{split}\boldsymbol{\tau}_{y}\odot\dot{\mathbf{y}}&=-\mathbf{y}+\mathbf{b}\odot\mathbf{z}+\left(\mathbf{1}-\mathbf{a}^{+}\right)\odot\left(\mathbf{W}_{r}\left(\sqrt{\mathbf{y}^{+}}-\sqrt{\mathbf{y}^{-}}\right)\right)\\ \boldsymbol{\tau}_{a}\odot\dot{\mathbf{a}}&=-\mathbf{a}+\mathbf{b}_{0}^{2}\odot\boldsymbol{\sigma}^{2}+\mathbf{W}\,\left(\left(\mathbf{y}^{+}+\mathbf{y}^{-}\right)\odot\mathbf{a}^{+2}\right)\end{split}$ (1)
@@ 5 Stability analysis for arbitrary recurrent weights
$\begin{split}\boldsymbol{\tau}_{y}\odot\dot{\mathbf{y}}&=-\mathbf{y}+\mathbf{b}\odot\mathbf{z}+\left(\mathbf{1}-\sqrt{\lfloor\mathbf{a}\rfloor}\right)\odot\left(\mathbf{W}_{r}\mathbf{y}\right)\\ \boldsymbol{\tau}_{a}\odot\dot{\mathbf{a}}&=-\mathbf{a}+\mathbf{b}_{0}^{2}\odot\boldsymbol{\sigma}^{2}+\mathbf{W}\,\left(\mathbf{y}^{2}\odot\lfloor\mathbf{a}\rfloor\right)\end{split}$ (15)
@@ 7 Discussion
The authors acknowledge valuable discussions with Flaviano Morone, Asit Pal, Mathias Casiulis, and Guanming Zhang. This work was supported by the National Institute of Health under award number R01EY035242. S.M. acknowledges the Simons Center for Computational Physical Chemistry for financial support. This work was supported in part through the NYU IT High-Performance Computing resources, services, and staff expertise.
@@ Appendix A Derivation of ORGaNICs
$\begin{split}\boldsymbol{\tau}_{y}\odot\dot{\mathbf{y}}&=-\mathbf{y}+\mathbf{b}\odot\mathbf{z}+\left(\mathbf{1}-\mathbf{a}^{+}\right)\odot\left(\mathbf{W}_{r}\left(\sqrt{\mathbf{y}^{+}}-\sqrt{\mathbf{y}^{-}}\right)\right)\\ \boldsymbol{\tau}_{a}\odot\dot{\mathbf{a}}&=-\mathbf{a}+\mathbf{b}_{0}^{2}\odot\boldsymbol{\sigma}^{2}+\mathbf{W}\,\left(\left(\mathbf{y}^{+}+\mathbf{y}^{-}\right)\odot\mathbf{a}^{+2}\right)\end{split}$ (34)
@@
$\begin{split}\boldsymbol{\tau}_{y}\odot\dot{\mathbf{y}}&=-\mathbf{y}+\mathbf{b}\odot\mathbf{z}+\left(\mathbf{1}-\mathbf{a}^{+}\right)\odot\left(\mathbf{W}_{r}\left(\left(\mathbf{y}^{+}\right)^{1/p}-\left(\mathbf{y}^{-}\right)^{1/p}\right)\right)\\ \boldsymbol{\tau}_{a}\odot\dot{\mathbf{a}}&=-\mathbf{a}+\frac{\mathbf{a}}{\left(\mathbf{a}^{+}\right)^{p}}\odot\left[\mathbf{b}_{0}^{p}\odot\boldsymbol{\sigma}^{p}+\mathbf{W}\left(\left(\mathbf{y}^{+}+\mathbf{y}^{-}\right)\odot\left(\mathbf{a}^{+}\right)^{p}\right)\right]\end{split}$ (35)
@@ Appendix F Analysis of the Rectified model
$\begin{split}\boldsymbol{\tau}_{y}\odot\dot{\mathbf{y}}&=-\mathbf{y}+\mathbf{b}\odot\mathbf{z}+\left(\mathbf{1}-\mathbf{a}^{+}\right)\odot\left\lfloor\mathbf{W}_{r}\mathbf{y}\right\rfloor\\ \boldsymbol{\tau}_{a}\odot\dot{\mathbf{a}}&=-\mathbf{a}+\mathbf{b}_{0}^{2}\odot\boldsymbol{\sigma}^{2}+\mathbf{W}\,\left(\mathbf{y}^{+}\odot\mathbf{a}^{+2}\right)\end{split}$ (87)
@@ F.3 Linear stability analysis of the two-dimensional model
$\begin{split}\tau_{y}\dot{y}&=-y+bz+\left(1-\sqrt{\lfloor a\rfloor}\right)\lfloor w_{r}y\rfloor\\ \tau_{a}\dot{a}&=-a+b_{0}^{2}\sigma^{2}+w\lfloor y\rfloor^{2}\lfloor a\rfloor\end{split}$ (102)
@@ Appendix G Iterative algorithm
$\begin{split}\boldsymbol{\tau}_{y}\odot\dot{\mathbf{y}}&=-\mathbf{y}+\mathbf{b}\odot\mathbf{z}+\left(\mathbf{1}-\sqrt{\lfloor\mathbf{a}\rfloor}\right)\odot\left(\mathbf{W}_{r}\mathbf{y}\right)\\ \boldsymbol{\tau}_{a}\odot\dot{\mathbf{a}}&=-\mathbf{a}+\mathbf{b}_{0}^{2}\odot\boldsymbol{\sigma}^{2}+\mathbf{W}\,\left(\mathbf{y}^{2}\odot\lfloor\mathbf{a}\rfloor\right)\end{split}$ (109)
@@ I.1 Static MNIST input
$\begin{split}\boldsymbol{\tau}_{y}\odot\dot{\mathbf{y}}&=-\mathbf{y}+f(\mathbf{W}_{bx}\mathbf{x})\odot\left(\mathbf{W}_{zx}\mathbf{x}\right)+\left(\mathbf{1}-\sqrt{\lfloor\mathbf{a}\rfloor}\right)\odot\left(\mathbf{W}_{r}\mathbf{y}\right)\\ \boldsymbol{\tau}_{a}\odot\dot{\mathbf{a}}&=-\mathbf{a}+\mathbf{b}_{0}^{2}\odot\boldsymbol{\sigma}^{2}+\mathbf{W}\,\left(\mathbf{y}^{2}\odot\lfloor\mathbf{a}\rfloor\right)\end{split}$ (134)
@@ I.2 Permuted and Unpermuted sequential MNIST
We performed a random split of 57,000 training samples and 3,000 validation samples and picked the model with the largest validation accuracy for testing. The unpermuted sequential MNIST task is defined as follows: for a given $28\times 28$ image, we flatten it to get a one-dimensional, 784 timestep-long input. Then these pixels are presented as an input ($\mathbf{x}_{i}$, one pixel at each time-step $i$) to the Euler discretized rectified ORGaNICs model (Eq. 87) with rectified input drive, given by the following equations,
$\begin{split}\mathbf{y}_{i+1}&=\mathbf{y}_{i}+\frac{\Delta t}{\boldsymbol{\tau}_{y}}\odot\left(-\mathbf{y}_{i}+\mathbf{b}_{i}\odot\left\lfloor\mathbf{W}_{zx}\mathbf{x}_{i}\right\rfloor+\left(\mathbf{1}-\mathbf{a}_{i}^{+}\right)\odot\left\lfloor\mathbf{W}_{r}\mathbf{y}_{i}\right\rfloor\right)\\ \mathbf{a}_{i+1}&=\mathbf{a}_{i}+\frac{\Delta t}{\boldsymbol{\tau}_{a}}\odot\left(-\mathbf{a}_{i}+\mathbf{b}_{0,i}^{2}\odot\boldsymbol{\sigma}^{2}+\mathbf{W}\,\left(\mathbf{y}_{i}^{+}\odot\mathbf{a}_{i}^{+2}\right)\right)\\ \mathbf{b}_{i+1}&=\mathbf{b}_{i}+\frac{\Delta t}{\boldsymbol{\tau}_{b}}\odot\left(-\mathbf{b}_{i}+f(\mathbf{W}_{bx}\mathbf{x}_{i}+\mathbf{W}_{by}\mathbf{y}_{i}+\mathbf{W}_{ba}\mathbf{a}_{i})\right)\\ \mathbf{b}_{0,i+1}&=\mathbf{b}_{0,i}+\frac{\Delta t}{\boldsymbol{\tau}_{b_{0}}}\odot\left(-\mathbf{b}_{0,i}+f(\mathbf{W}_{b_{0}x}\mathbf{x}_{i}+\mathbf{W}_{b_{0}y}\mathbf{y}_{i}+\mathbf{W}_{b_{0}a}\mathbf{a}_{i})\right)\end{split}$ (135)
When we are done presenting the pixels we use the last hidden state, i.e., $\mathbf{y}_{784}$, to make the predictions. To make this more challenging we also train ORGaNICs on permuted sMNIST where we first permute the pixels of all the images in some random order and the rest of the task is the same. Instead of parametrizing $\boldsymbol{\tau}$, we parametrize ${\Delta t}/{\boldsymbol{\tau}_{y}}=0.05*f(\mathbf{p}_{y})$, ${\Delta t}/{\boldsymbol{\tau}_{a}}=0.01*f(\mathbf{p}_{a})$ and ${\Delta t}/{\boldsymbol{\tau}_{b}}=0.1*f(\mathbf{p}_{b})$ and ${\Delta t}/{\boldsymbol{\tau}_{b_{0}}}=0.1*f(\mathbf{p}_{b_{0}})$, so we can control the dimensionless relative time constants. In practice, we find it is better to make the $\mathbf{a}$ neurons sluggish compared to $\mathbf{y}$. This is based on the intuition given by the two-dimensional phase portrait for different relative time constants Fig. 3. All the parameters (including $\mathbf{W}_{r}$) are unconstrained for this task with initialization specified in Table 6. Since ORGaNICs are stable, we did not need to use gradient clipping for training, which is commonly used for LSTMs. Additionally, we train the model using a StepLR learning rate scheduler with parameters given in Table 7. The model is trained using backpropagation through time (BPTT) and takes approximately 30 hours to fully train.